OpenStudy (anonymous):

I am having trouble understanding how to solve polynomials for 0 when there are more then one fraction in the equation. Is there a trick I am missing?

OpenStudy (anonymous):

Do you have an example of a question?

OpenStudy (anonymous):

I do, \[(z ^{2})/6-(2z)/2-3=0\]

OpenStudy (bahrom7893):

Nice picture joemath!

OpenStudy (anonymous):

I know what do if there is only one fraction in the polynomial but I am totally lost with 2. I missed class :/

OpenStudy (anonymous):

lol thanks :P for this problem: \[\frac{z^2}{6}-\frac{2z}{2} - 3 = 0\] what you want to do is look at the denominators, and think of one number, that when you multiply the equation by this number, will get rid of them. It turns out to be the LCM (least common multiple) of the denominators.

OpenStudy (anonymous):

So I pull out two and then leave the 3 under the first?

OpenStudy (anonymous):

if the equation I typed in my last response correct?

OpenStudy (anonymous):

er, not if, was. was the equation correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, so the denominators are 2, and 6. Their LCM turns out to also be 6. So multiplying everything by we get: \[6*\frac{z^2}{6}+6*\frac{2z}{2}-6*3= 6*0 \Rightarrow z^2+3*2z-18 = 0 \Rightarrow z^2+6z-18 = 0\] No More fractions :)

OpenStudy (anonymous):

Oh... Thanks.. I see what I was doing vs what needs to be done. Thanks again. Back to homework.